OpenStudy (anonymous):

Imagine all 7 billion people on earth wanted to gather in one place. Let's assume we allow one square meter per person, and that we all gather inside a circle. How many kilmeters are in the radius of the circle? Round your answer to the neared whole kilometer. Use pi=3.14 in your calculations. Imagine all 7 billion people on earth wanted to gather in one place. Let's assume we allow one square meter per person, and that we all gather inside a circle. How many kilmeters are in the radius of the circle? Round your answer to the neared whole kilometer. Use pi=3.14 in your calculations. @Mathematics

OpenStudy (anonymous):

what is 7 billion ?

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

what omg ?

OpenStudy (anonymous):

come on please ans. im runing out of time.

OpenStudy (anonymous):

how many zeros in 7 billion?

OpenStudy (anonymous):

9 zeroes

OpenStudy (anonymous):

Good...

OpenStudy (anonymous):

i can't give medals y?

OpenStudy (anonymous):

No idea ..

OpenStudy (agreene):

I'd like to point out that the current, most accurate measurement of world population is from 2010 and it is 6,840,507,000 I don't appreciate that much rounding.

OpenStudy (anonymous):

\[ 10 \sqrt{\frac{70}{\pi }} \] Km

OpenStudy (anonymous):

@agreene:It is not correct, the number is 6,840,507,001 they forgot to count me :P